OpenStudy (anonymous):

Sketch the region enclosed by the given curves. y = 5 cos 2x, y = 5 − 5 cos 2x, 0 ≤ x ≤ π/2 what its area

OpenStudy (anonymous):

\[\int\limits_{0}^{2\theta}\int\limits_{0}^{5-5\cos 2x}rdrd \theta\]

OpenStudy (anonymous):

whats r?

OpenStudy (anonymous):

What level are you? Have you done double integrals?